If the tens digit x and the unit digit y of a positive integer n are reversed, the resulting integer 9 is more than n, what is y in terms of x?
(A)10-x
(B)9-x
(C)x+9
(D)x-1
(E)x+1
The answer is E, x+1 but I am not getting it. Could you please explain.

Thanks

Answer by Edwin McCravy(20086)

If the tens digit x and the unit digit y of a positive integer n are reversed, the resulting integer 9 is more than n, what is y in terms of x?

The given positive integer n = 10x + y

The resulting positive integer when the integers are reversed is 10y + x

(10y + x) = (10x + y) + 9

  10y + x = 10x + y + 9

Add -y - x to both sides

   9y = 9x + 9

    y = x + 1

Edwin